OpenStudy (anonymous):

The circumference of the Earth is approximately 25,000 miles. Find the diameter of the Earth. A. 7,957.8 miles B. 7,958.7 miles C. 7,958.6 miles D. 7,957.7 miles

OpenStudy (apoorvk):

7954.54.. this is what you get when you take pi=22/7

OpenStudy (anonymous):

I got 7961.7?

OpenStudy (apoorvk):

7961.35... whta you get with pi=3.14

OpenStudy (anonymous):

Than what the answer?

OpenStudy (apoorvk):

well this is tricky... cause. taking pi=22/7 should give an accurate answer, cause approximations aren't involved. yet your options seem fishy. cause we are landing quite far from it. can you recheck the options provided once?

OpenStudy (apoorvk):

hmm. well seems to be a problem. i ll check that out. can you recheck the calculations yourself using a calculator if you can, and reply to verify?

OpenStudy (anonymous):

D = C/pi D is the diameter C is the Circumference D=25000/pi which is 7957.747. if you round up it will be 7957.75 and since it is asking for only the tenths place, you round up once more for 7957.8 miles. The answer is A.

OpenStudy (anonymous):

My bad, you aren't supposed to look at the digits past the one you are rounding. So 7957.74 is all you need to look at since you are rounding to the tenths place. Therefore, the answer is 7957.7 letter D. Sorry about that last one!